A 1000 kg space craft launched from the Earth is now passing Pluto at a distance of 7.5 Billion kilometers from the Sun, but its engines will shut down because its fuel has run out. What must be its speed in order to escape the gravitational pull of the sun? (Use Msun = 1.989 x 10^30 kg)
Answer #1

Let speed be v near Pluto ,

For escaping, KE near Pluto + PE near Pluto = KE at infinite distance + PE at infinite distance

0.5mv^2 - GmM/r = 0 + 0

0.5*1000*v^2 -6.67e-11*1000*1.989e30/7.5e12 = 0

v = sqrt(2*6.67e-11*1.989e30/7.5e12)

= 5948 m/s answer